Beer costs R10 per bottle and cigarettes cost R40 per packet. Suppose Vongani buys 8 bottles of beer and 3 packets of cigarettes and that his marginal utility from beer is then 20 utils and that from cigarettes 35 utils


1
Expert's answer
2021-09-06T06:59:49-0400

In equilibrium:

"\frac{MU1} {P1} = \frac{MU2} {P2}"


and


"P1 \times Q1 + P2 \times Q2 = I"

In our case "20\/10>35\/40" , so John should consume more beer and less cigarettes to maximize his utility.